OpenStudy (anonymous):

What is a good estimate for the square root of 51?

OpenStudy (traylea):

Estimate 51 to the nearest 10 and get it from there?

OpenStudy (anonymous):

10

OpenStudy (welshfella):

Hint 51 is very close to 49

OpenStudy (welshfella):

- and square root of 49 is what?

OpenStudy (anonymous):

7

OpenStudy (welshfella):

so square root of 51 will be a little bigger than 7

OpenStudy (welshfella):

one pretty good method of doing this is to take the next perfect square after 49 which is 64 Then subtract them: 64-49=15 . Now 51-49 = 2 . Now we add 2/15 to 7 . 2/15 = 0.13. so we get 7 + 0.13 = 7.13 as the estimate.

OpenStudy (welshfella):

If you square 7.13 on your calcukator you get 50.84 Pretty close to 51!

OpenStudy (mathstudent55):

This method is called linear interpolation.